Operations Report – May 25th

Crew 195 Operations Report 25.05.18

SOL: 5

Name of person filing report: Dana Levin

Non-nominal systems: Continuing run Diesel generator at night. No service rep for Solar yet.

Notes on non-nominal systems: n/a

Generator (hours run): Overnight approximately 8 hours, on solar until 1600, now back on generator

Solar— SOC 10%

Diesel – Over 50%

Propane – 50 PSI

ATV Fuel (Ethanol free gasoline) – 1 gallon.

Water (trailer) – 485 gallons.

Water (static) – Approximately 90

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1351243

Toilet tank emptied: No
Date last emptied: 5.22.18

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda and 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 1 gallon

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: Honda and ATV 300 out of service.

Deimos rover used: yes
Hours: 1 hour
Beginning charge: 85%
Ending charge: 80 %
Currently charging: Yes

Curiosity rover used: Yes

Hours: 1 hour
Beginning charge: 99%
Ending charge: 85 %
Currently charging: Yes

Spirit rover used: no

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Working well

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

Operations Report – May 23rd

Crew 195 Operations Report 23.05.18

SOL: 3

Name of person filing report: Chris Davis

Non-nominal systems: Continuing run Diesel generator at night. No service rep for Solar yet. Septic serviced yesterday

Notes on non-nominal systems:

Generator (hours run): Overnight approximately 10 plus 4 hours today per Shannon

Solar— SOC 10%

Diesel – Over 50%

Propane – 51 PSI

ATV Fuel (Ethanol free gasoline) – 4 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 90

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1351235

Toilet tank emptied: No
Date last emptied: 5.22.18

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda and 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 1 gallon

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: Honda and ATV 300 out of service.

Deimos rover used: yes
Hours: N/A
Beginning charge: 100%
Ending charge: 69 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes

Beginning charge: 100 %
Ending charge: 71%
Currently charging: Yes

Opportunity rover used: Yes
Beginning charge: 96%
Ending charge: 64%
Currently charging: NA

Curiosity rover used: No
Beginning charge: 100 %
Ending charge: 71%
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Working well

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

Operations Report – May 22nd

[tittle Operations Report – May 22]

Crew 195 Operations Report 22.05.18

SOL: 2

Name of person filing report: Chris Davis

Non-nominal systems: Continuing run Diesel generator at night. No service rep for Solar yet. Septic serviced today.

Notes on non-nominal systems:

Generator (hours run): Overnight approximately 10 plus 4 hours today per Shannon

Solar— SOC 0%

Diesel – Over 50%

Propane – 51 PSI

ATV Fuel (Ethanol free gasoline) – 4 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 150

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1351176

Toilet tank emptied: Yes
Date last emptied: 5.22.18

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda and 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 1 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: Honda and ATV 300 out of service.

Deimos rover used: yes
Hours: N/A
Beginning charge: 100%
Ending charge: 74 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes

Beginning charge: 100 %
Ending charge: 76%
Currently charging: Yes

Opportunity rover used: No
Beginning charge: NA
Ending charge: NA
Currently charging: NA

Curiosity rover used: No
Beginning charge: 100 %
Ending charge: 71%
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Working well

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

Operations Report – May 20th

Crew 195 Operations Report 20.05.18

SOL: 0

Name of person filing report: Chris Davis

Non-nominal systems: Continuing run Diesel generator at night. No service rep for Solar yet.

Notes on non-nominal systems:

Generator (hours run): Overnight approximately 10 plus 4 hours today per Shannon

Solar— SOC 52%

Diesel – Over 50%

Propane – 51 PSI

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 250

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1351130

Toilet tank emptied: No
Date last emptied: Previous Crew Sol 5

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda and 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 3 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: Honda and ATV 300 out of service.

Deimos rover used: yes
Hours: N/A
Beginning charge: 96%
Ending charge: 82 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes

Hours: 35.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 92 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Working well

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

Operations Report – May 18th

Crew 194 Operations Report 18.05.18

SOL: 5

Name of person filing report: Ben Easter

Non-nominal systems: Continuing to work on power system. Generator run during day today per Shannon’s instructions.

Notes on non-nominal systems:

Generator (hours run): Overnight approximately 10 plus 4 hours today per Shannon

Solar— SOC 52%

Diesel – Over 50%

Propane – psi. Not checked

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 250

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 135115

Toilet tank emptied: Yes
Date last emptied: Today

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 3 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: We recovered ATV 300.

Deimos rover used: No
Hours: N/A
Beginning charge: 100%
Ending charge: XX %
Currently charging: Yes or No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes

Hours: 35.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 94 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: XX %
Currently charging: Yes or No

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Working well

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

Operations Report – 17 May 2018

SOL: 4

Name of person filing report: Ben Easter

Non-nominal systems: Power per prior documentation. We attempted to turn on the charger today per Shannon’s recommendations, but it would not power on. Pending service visit in a few days.

Notes on non-nominal systems: Continued issues as already described

Generator (hours run): Overnight approximately 10

Solar— SOC 0%

Diesel – Over 50%

Propane – psi. Not checked

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 50

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 135037

Toilet tank emptied: Yes
Date last emptied: Today

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 3 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: ATV 300 battery died approximately 500 meters from Hab. Have removed battery and are charging.

Deimos rover used: No
Hours: N/A
Beginning charge: 100%
Ending charge: XX %
Currently charging: Yes or No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes

Hours: 35.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 94 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: XX %
Currently charging: Yes or No

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Slow today

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

Operations Report – May 16th

Crew 194 Operations Report 16.05.18

SOL: 3

Name of person filing report: Ben Easter

Non-nominal systems: Power per prior documentation

Notes on non-nominal systems: Continued issues as already described

Generator (hours run): Overnight approximately 10

Solar— SOC 0%

Diesel – Over 50%

Propane – psi. Not checked

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 75%

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 134955

Toilet tank emptied: No
Date last emptied: By director prior to our arrival

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 3 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs:

Deimos rover used: No
Hours: N/A
Beginning charge: 100%
Ending charge: XX %
Currently charging: Yes or No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes

Hours: 35.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 94 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: XX %
Currently charging: Yes or No

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: working well

Summary of suits and radios: working well

Summary of Hab operations: Intermittent gas odor with turning on water heater.

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

Operations Report – May 15th

Crew 194 Operations Report 15.05.18

SOL: 2

Name of person filing report: Ben Easter

Non-nominal systems: Power per prior documentation

Notes on non-nominal systems: Continued issues as already described

Generator (hours run): Overnight approximately 10

Solar— SOC 56%

Diesel – 60%

Propane – psi. Not checked

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 75%

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 134890

Toilet tank emptied: Yes
Date last emptied: By director prior to our arrival

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 3 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs:

Deimos rover used: No
Hours: N/A
Beginning charge: 100%
Ending charge: XX %
Currently charging: Yes or No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 35.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 94 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: XX %
Currently charging: Yes or No

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: working well

Summary of suits and radios: working well

Summary of Hab operations: no concerns

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: None.

Operations Report – May 15th

Crew 194 Operations Report 15.05.18

SOL: 2

Name of person filing report: Ben Easter

Non-nominal systems: Power per prior documentation

Notes on non-nominal systems: Continued issues as already described

Generator (hours run): Overnight approximately 10

Solar— SOC 56%

Diesel – 60%

Propane – psi. Not checked

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 75%

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 134890

Toilet tank emptied: Yes
Date last emptied: By director prior to our arrival

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 3 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs:

Deimos rover used: No
Hours: N/A
Beginning charge: 100%
Ending charge: XX %
Currently charging: Yes or No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 35.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 94 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: XX %
Currently charging: Yes or No

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: working well

Summary of suits and radios: working well

Summary of Hab operations: no concerns

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: One crew member experienced a medical incident during the EVA related to heat exhaustion while exerting in the heat/sun. The EVA was aborted, and the crew returned to the Hab. The crew member was treated appropriately for heat illness.

Questions, concerns and requests to Mission Support: None.